OpenStudy (anonymous):

Two numbers have a sum of 18. If the small number doubled, and then subtracted from the larger number, the difference is 9. X represents the smaller number, Y represents the biggest number. Write an equation to model the difference of the numbers.

OpenStudy (anonymous):

Use the equations \[18 = x +y\] and \[9 = y - 2x \]